Shift report 20010901 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sat Sep 01 2001 - 00:08:10 EDT

  • Next message: Apache: "Shift report 20010901 00:00-08:00"

    ------------------------------------------------------------
    Date: Sat Sep  1 00:08:09 2001, Shift: 16:00-24:00
    Supervisor: ZM, with: -
    Log book#8 pages 50 - 51
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam stored and cogged at  19:50.
       Goog beam up to end of the shift. 
       started at B=13.6x10^9, Y=18.1x10^9 and decreased to
       B=10.8x10^9, Y=16.0x10^9 
    2. Two runs (4530 and 4531) with the same setting recorded
       (About 0.5M trigger #1 event)
       MRS at 40 deg., no magnetic field, FS at 4 deg, polarity B
       1/5 mag. field.
    3. No detector problems.                    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sat Sep 01 2001 - 00:08:50 EDT